# ChatEMG: Synthetic Data Generation to Control a Robotic Hand Orthosis for Stroke

### v1

@@ VI-A Intent Inferral Performance
S1 S2 S3 S4 S5 Avg.
LDA Self $0.37\pm 0.15$ $0.54\pm 0.07$ $0.46\pm 0.10$ $0.64\pm 0.22$ $0.36\pm 0.07$ $0.48$
Fine-tune $0.34\pm 0.12$ $0.64\pm 0.10$ $0.54\pm 0.05$ $0.71\pm 0.05$ $0.48\pm 0.15$ $0.54$
RF Self $0.52\pm 0.07$ $0.72\pm 0.08$ $0.64_{\pm 0.10}$ $0.90\pm 0.08$ $0.77\pm 0.06$ $0.71$
Fine-tune $0.53\pm 0.11$ $0.68\pm 0.06$ $0.63\pm 0.04$ $0.89\pm 0.08$ $0.64\pm 0.17$ $0.67$
Transformer Self $0.55\pm 0.05$ $0.64\pm 0.05$ $0.58\pm 0.10$ $0.84\pm 0.06$ $0.73\pm 0.07$ $0.67$
Fine-tune $0.57\pm 0.06$ $0.64\pm 0.05$ $0.70\pm 0.03$ $0.88_{\pm 0.04}$ $0.68\pm 0.06$ $0.69$
@@
LDA Self $0.40\pm 0.16$ $0.50\pm 0.09$ $0.52\pm 0.07$ $0.50\pm 0.13$ $0.81\pm 0.05$ $0.54$
Fine-tune $0.64\pm 0.04$ $0.49\pm 0.07$ $0.65\pm 0.05$ $0.65\pm 0.02$ $0.64\pm 0.13$ $0.61$
RF Self $0.55\pm 0.08$ $0.57\pm 0.07$ $0.77\pm 0.11$ $0.82\pm 0.13$ $0.77\pm 0.06$ $0.69$
Fine-tune $0.58\pm 0.06$ $0.62\pm 0.02$ $0.74\pm 0.16$ $0.72\pm 0.16$ $0.78\pm 0.05$ $0.69$
Transformer Self $0.66\pm 0.06$ $0.53\pm 0.05$ $0.77\pm 0.09$ $0.67\pm 0.15$ $0.72\pm 0.15$ $0.67$
Fine-tune $0.72\pm 0.06$ $0.57\pm 0.12$ $0.78\pm 0.12$ $0.79\pm 0.09$ $0.79\pm 0.08$ 0.73
@@
LDA Self $0.33\pm 0.06$ $0.61\pm 0.11$ $0.51\pm 0.11$ $0.52\pm 0.26$ $0.65\pm 0.18$ $0.51$
Fine-tune $0.60\pm 0.08$ $0.62\pm 0.10$ $0.49\pm 0.06$ $0.74\pm 0.03$ $0.40\pm 0.17$ $0.56$
@@
RF Self $0.60\pm 0.01$ $0.64\pm 0.02$ $0.59\pm 0.02$ $0.80\pm 0.08$ $0.71\pm 0.07$ $0.66$
Fine-tune $0.54\pm 0.02$ $0.65\pm 0.02$ $0.55\pm 0.12$ $0.78\pm 0.07$ $0.75\pm 0.13$ $0.65$
@@
Transformer Self $0.57\pm 0.05$ $0.63\pm 0.03$ $0.63\pm 0.06$ $0.72\pm 0.04$ $0.65\pm 0.16$ $0.64$
Fine-tune $0.56\pm 0.08$ $0.64\pm 0.02$ $0.60\pm 0.10$ $0.68\pm 0.08$ $0.69\pm 0.04$ $0.64$
@@
TABLE III: Subject adaptation experiment results.
The results of condition, session, and subject adaptation experiments are shown in Table III and  III, respectively. For each subject, we evaluate three intent inferral recordings, and we present the average accuracy and one standard deviation. ChatEMG is able to improve the average intent inferral accuracy across five subjects for all three classifiers under all three assessment scenarios. This shows that ChatEMG can successfully generalize to a new condition, session, or subject despite not seeing them in its training recordings.
@@ VI-B Synthetic Sample Visualization and Analysis
We show generated samples for lower-functioning subjects S1, S2, and S3 in Fig. 5. The first 150-step (1.5s) EMG signals are the sampled prompt from the limited dataset and are identical between real and synthetic data. For the synthetic samples, the last 106-step (1.06s) EMG signals are generated by ChatEMG. When presented with both plots in parallel, without knowing in advance, it is very challenging to identify the synthetic one. There is no significant transition in signals at 1.5s when the sequence switches from real to synthetic. This shows that ChatEMG can capture the characteristics of EMG signals, such as the amplitude, frequency, fluctuation pattern, etc. ChatEMG can also maintain the relative position of different channels very well.
More importantly, ChatEMG not only learns to babble EMG signals by following the previous trends with the correct EMG signal characteristics, but it also learns to reproduce common trends that do not show up in the prompt at all. For example, in the prompt of S1’s open sample, there is no indication of the green channel (emg2) going up, but the generated sequence shows such a trend, which turns out to be correct.
We further investigate the generated samples of subject S4 by using t-SNE [25] to visualize them in a low-dimensional space, as shown in Fig. 6. For each class, we generate 1000 synthetic samples using the support set and then randomly select 100 samples from the query set of the same recording. We embed each channel of the 256-step EMG sequence separately into a 2D space, and visualize that of the first two channels. We have two observations from this plot. (1) The synthetic samples of different intents are very separable from each other, meaning that ChatEMG captures distinct patterns of different intents. (2) The embedding space of the generated samples almost always covers the real samples from the query set. This shows that ChatEMG captures the distribution of the test samples correctly. Thus, adding these synthetic samples to the limited support set can improve the intent inferral accuracy on the query set.
@@ VII Conclusion
We propose ChatEMG, an autoregressive generative model that can generate synthetic EMG signals conditioned on prompts. Instead of collecting a large labeled dataset from the new condition, session, or subject, ChatEMG allows us to collect only a very small dataset and expand it with synthetic samples. ChatEMG learns the broad behavior of forearm EMG signals from a vast corpus of previous data while remaining context-specific via prompting. We show that these synthetic samples are classifier-agnostic and can improve the intent inferral accuracy of different types of classifiers. We are the first to deploy an intent classifier trained partially on synthetic data on a hand orthosis to help an unseen stroke subject complete pick-and-place tasks, showing that the improvement in classification accuracy can lead to improvement in meaningful functional tasks.

### v3

@@ VI-A Intent Inferral Performance
$p$-value w/
ChatEMG
LDA Self $0.37\pm 0.15$ $0.54\pm 0.07$ $0.46\pm 0.10$ $0.64\pm 0.22$ $0.36\pm 0.07$ $0.48$ $\mathbf{3\mathrm{\textbf{e}}{-4}}$
Fine-tune $0.34\pm 0.12$ $0.64\pm 0.10$ $0.54\pm 0.05$ $0.71\pm 0.05$ $0.48\pm 0.15$ $0.54$ $\mathbf{4\mathrm{\textbf{e}}{-4}}$
RF Self $0.52\pm 0.07$ $0.72\pm 0.08$ $0.64\pm 0.10$ $0.90\pm 0.08$ $0.77\pm 0.06$ $0.71$ $1\mathrm{e}{-1}$
Fine-tune $0.53\pm 0.11$ $0.68\pm 0.06$ $0.63\pm 0.04$ $0.89\pm 0.08$ $0.64\pm 0.17$ $0.67$ $\mathbf{5\mathrm{\textbf{e}}{-2}}$
Transformer Self $0.55\pm 0.05$ $0.64\pm 0.05$ $0.58\pm 0.10$ $0.84\pm 0.06$ $0.73\pm 0.07$ $0.67$ $\mathbf{3\mathrm{\textbf{e}}{-2}}$
Fine-tune $0.57\pm 0.06$ $0.64\pm 0.05$ $0.70\pm 0.03$ $0.88\pm 0.04$ $0.68\pm 0.06$ $0.69$ $2\mathrm{e}{-1}$
@@
$p$-value w/
ChatEMG
LDA Self $0.40\pm 0.16$ $0.50\pm 0.09$ $0.52\pm 0.07$ $0.50\pm 0.13$ $0.81\pm 0.05$ $0.54$ $\mathbf{6\mathrm{\textbf{e}}{-5}}$
Fine-tune $0.64\pm 0.04$ $0.49\pm 0.07$ $0.65\pm 0.05$ $0.65\pm 0.02$ $0.64\pm 0.13$ $0.61$ $\mathbf{1\mathrm{\textbf{e}}{-2}}$
RF Self $0.55\pm 0.08$ $0.57\pm 0.07$ $0.77\pm 0.11$ $0.82\pm 0.13$ $0.77\pm 0.06$ $0.69$ $1\mathrm{e}{-1}$
Fine-tune $0.58\pm 0.06$ $0.62\pm 0.02$ $0.74\pm 0.16$ $0.72\pm 0.16$ $0.78\pm 0.05$ $0.69$ $2\mathrm{e}{-1}$
Transformer Self $0.66\pm 0.06$ $0.53\pm 0.05$ $0.77\pm 0.09$ $0.67\pm 0.15$ $0.72\pm 0.15$ $0.67$ $\mathbf{2\mathrm{\textbf{e}}{-2}}$
Fine-tune $0.72\pm 0.06$ $0.57\pm 0.12$ $0.78\pm 0.12$ $0.79\pm 0.09$ $0.79\pm 0.08$ 0.73 $5\mathrm{e}{-1}$
TABLE III: Subject adaptation experiment results.
S1 S2 S3 S4 S5 Average
$p$-value w/
ChatEMG
LDA Self $0.33\pm 0.06$ $0.61\pm 0.11$ $0.51\pm 0.11$ $0.52\pm 0.26$ $0.65\pm 0.18$ $0.51$ $\mathbf{1\mathrm{\textbf{e}}{-2}}$
Fine-tune $0.60\pm 0.08$ $0.62\pm 0.10$ $0.49\pm 0.06$ $0.74\pm 0.03$ $0.40\pm 0.17$ $0.56$ $3\mathrm{e}{-1}$
@@
RF Self $0.60\pm 0.01$ $0.64\pm 0.02$ $0.59\pm 0.02$ $0.80\pm 0.08$ $0.71\pm 0.07$ $0.66$ $3\mathrm{e}{-1}$
Fine-tune $0.54\pm 0.02$ $0.65\pm 0.02$ $0.55\pm 0.12$ $0.78\pm 0.07$ $0.75\pm 0.13$ $0.65$ $\mathbf{5\mathrm{\textbf{e}}{-2}}$
@@
Transformer Self $0.57\pm 0.05$ $0.63\pm 0.03$ $0.63\pm 0.06$ $0.72\pm 0.04$ $0.65\pm 0.16$ $0.64$ $\mathbf{1\mathrm{\textbf{e}}{-2}}$
Fine-tune $0.56\pm 0.08$ $0.64\pm 0.02$ $0.60\pm 0.10$ $0.68\pm 0.08$ $0.69\pm 0.04$ $0.64$ $\mathbf{4\mathrm{\textbf{e}}{-2}}$
@@
The results are shown in Table III and III. For each subject, we evaluate three intent inferral recordings, and we present the average accuracy and one standard deviation. ChatEMG is able to improve the average intent inferral accuracy across five subjects for all classifiers under all assessment scenarios. This shows that ChatEMG can successfully generalize to a new condition, session, or subject despite not seeing them in its training recordings.
Despite that the improvement in intent inferral accuracy is consistent across 17/18 comparisons between ChatEMG and Fine-tune/Self, we further investigate the statistical significance of such improvement by performing a one-sided Wilcoxon rank-sum test on the results aggregated across all subjects (three intent inferral recordings per subject). We choose a non-parametric statistical test because we do not assume an underlying normal distribution. We report the computed $p$-values for pairwise differences between our ChatEMG and the other methods. With a commonly used hypothesis threshold of $\alpha=5\mathrm{e}{-2}$, 11/17 improvements ($p$-values in bold) are statistically significant.
@@ VI-B Synthetic Sample Visualization and Analysis
We show generated samples for lower-functioning subjects S1, S2, and S3 in Fig. 5. The first 150 steps (1.5s) are the sampled prompts from the limited dataset and are identical between real and synthetic data. For the synthetic samples, ChatEMG generates the last 106 steps (1.06s). When presented with both plots in parallel, without knowing in advance, it is very challenging to identify the synthetic one. There is no significant transition at 1.5s when the sequence switches from real to synthetic. This shows that ChatEMG can capture the characteristics of EMG signals, such as the amplitude, frequency, fluctuation pattern, etc. ChatEMG can also maintain the relative position of different channels very well. We compute the normalized root mean squared error (NRMSE) between the synthetic and real samples across all subjects and intents. The NRMSE for close, open, and relax intents across all subjects are 6%, 5%, and 3%, respectively.
More importantly, ChatEMG not only learns to babble EMG signals by following the previous trends, but it also learns to reproduce common trends that do not show up in the prompt at all. For example, in the prompt of S1’s open sample, there is no indication of the green channel (emg2) going up, but the generated sequence shows such a trend, which turns out to be correct.
We further visualize the generated samples of subject S4 in a low-dimensional space using t-SNE [26], shown in Fig. 6. For each class, we generate 1000 synthetic samples using the support set and then randomly select 100 samples from the query set of the same recording. We embed each channel of the 256-step EMG sequence separately into a 2D space, and visualize that of the first two channels. We observe that: (1) The synthetic samples of different intents are very separable from each other, meaning that ChatEMG captures distinct patterns of different intents. (2) The embedding space of the generated samples almost always covers the real samples from the query set. This shows that ChatEMG captures the distribution of the test samples correctly. Thus, adding these synthetic samples to the limited support set can improve the intent inferral accuracy on the query set.
@@ VII Conclusion
We propose ChatEMG, an autoregressive generative model that can generate synthetic EMG signals conditioned on prompts. ChatEMG allows us to collect only a very small dataset from the new condition/session/subject and expand it with synthetic samples. ChatEMG learns the broad behavior of forearm EMG signals from a vast corpus of previous data while remaining context-specific via prompting. We show that these synthetic samples are classifier-agnostic and can improve the intent inferral accuracy of different types of classifiers. We are the first to deploy an intent classifier trained partially on synthetic data on a hand orthosis to help an unseen stroke subject complete pick-and-place tasks, showing that the improvement in classification accuracy can lead to improvement in meaningful functional tasks.